# ![[Uncaptioned image]](extracted/6365550/Fig/teaser.png) Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution

### v2

@@ Abstract.
Chain-of-Thought (CoT) prompting enhances the reasoning of large language models (LLMs) by decomposing problems into sequential steps, mimicking human logic and reducing errors. However, complex tasks with vast solution spaces and vague constraints often exceed the capacity of a single reasoning chain. Inspired by Minimal Free Resolution (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a novel framework that extends CoT by introducing auxiliary, interrelated reasoning paths. SoT captures deeper logical dependencies, enabling more robust and structured problem-solving. MFR decomposes a module into a sequence of free modules with minimal rank, providing a structured analytical approach to complex systems. This method introduces the concepts of ”Module”, ”Betti numbers”,”Freeness”, ”Mapping”, ”Exactness” and ”Minimality”, enabling the systematic decomposition of the original complex problem into logically complete minimal subproblems while preserving key problem features and reducing reasoning length. We tested SoT across diverse datasets (e.g., GSM8K, MATH) and models (e.g., GPT-4o-mini, Qwen2.5), achieving inference accuracy that matches or surpasses mainstream CoTs standards. Additionally, by aligning the sampling process with algebraic constraints, our approach enhances the scalability of inference time in LLMs, ensuring both transparent reasoning and high performance. Our code will be publicly available at https://github.com/dlMARiA/Syzygy-of-thoughts
1. Introduction
@@
The rapid advancements in natural language processing (Vaswani et al., 2017; Devlin et al., 2019; Radford et al., 2018; Brown et al., 2020) have demonstrated outstanding reasoning capabilities across various tasks (Zhao et al., 2023b; Min et al., 2023). In recent years, the Chain of Thought (CoT) prompting method (Wei et al., 2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems (Chen et al., 2025; Qu et al., 2025; Sui et al., 2025; Wei et al., 2025; Xia et al., 2024; Liu et al., 2025). The core advantage of the CoT framework lies in its ability to decompose problems into a series of explicit and easily interpretable intermediate steps, thereby improving the transparency and explainability of the reasoning process. This explicit stepwise reasoning not only facilitates in-depth analysis and debugging of model outputs but also provides a clear logical pathway for solving complex problems.
@@
However, despite its success in many tasks, CoT exhibits inherent limitations when dealing with high-dimensional, nonlinear, and abstract logical problems (Wang et al., 2022a; Shi et al., 2023; Jin et al., 2024). First, traditional CoT methods often rely on fixed or heuristic decomposition strategies, which struggle to capture essential details and intrinsic connections in multidimensional problems (Wang et al., 2022a). Second, due to the lack of systematic structural planning in intermediate steps, the reasoning process tends to include redundant computations and repetitive paths, leading to increased computational resource consumption and inference latency (Jin et al., 2024). Although existing improvements have alleviated these issues to some extent, introducing techniques such as multi-path exploration, voting mechanisms, and hierarchical decomposition strategies to enhance robustness (Yao et al., 2023a; Besta et al., 2024; Wang et al., 2022b; Zhou et al., 2022; Yao et al., 2023b), however, they have yet to fully address logical incompleteness and the inadequate decomposition of high-dimensional information. As a result, when handling highly complex reasoning tasks, model robustness and accuracy still face significant challenges.
@@
To address these challenges, we propose a novel reasoning framework—Syzygy of Thoughts (SoT). This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra (Eisenbud, 2013; Rossi and Sharifan, 2009; Botbol et al., 2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences. The incorporation of MFR theory ensures a more rigorous and efficient problem decomposition process while preserving the core characteristics of the problem, effectively reducing redundant computations and logical inconsistencies. Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as ”module”, ”Betti numbers”, ”freeness”, ”mapping”, ”exactness” and ”minimality” to deeply explore and systematically express the intrinsic structure of complex problems. This approach provides a novel theoretical perspective for analyzing high-dimensional and multivariable problems.
@@
Finally, the SoT framework exhibits multiple advantages, including an intuitive and transparent reasoning path and improved interpretability of intermediate information. The main contributions of this work are as follows:
Pioneering the integration of MFR into LLM reasoning, enabling structured decomposition of complex problems.
Developing a novel framework that builds on CoT’s strengths while substantially improving reasoning clarity and structural efficiency.
Providing theoretical insights and empirical evidence that SoT significantly enhances performance in high-dimensional reasoning tasks.
2. Related Work
2.1. Advanced Reasoning Techniques in LLMs
The reasoning capabilities of LLMs have advanced significantly with breakthroughs in prompting paradigms, notably CoT prompting proposed by Wei et al.(Wei et al., 2022). CoT decomposes complex problems into stepwise solutions, mimicking human reasoning, yet struggles with multidimensional or nonlinear tasks like abstract logic and advanced mathematics. To address these limitations, variants such as Zero-shot-CoT(Kojima et al., 2022), Self-Consistency CoT (Wang et al., 2022b), and Auto-CoT (Zhang et al., 2022) have emerged, alongside verification-based approaches like VerifyCoT (Zhao et al., 2023a) and CoF-CoT (Nguyen et al., 2023), yielding measurable improvements.
Structured prompting strategies further enhance reasoning. Least-to-Most Prompting (LtM)(Zhou et al., 2022) breaks tasks into manageable subproblems, reducing single-step errors. Programmatic methods, including Program of Thought (PoT)(Chen et al., 2022), Chain of Code (CoC)(Li et al., 2023), and Buffer of Thought (BoT)(Yang et al., 2025), integrate discrete variables and procedural steps. Building on these, Algorithm of Thought (AoT) (Sel et al., 2023) iteratively synthesizes and refines solutions, boosting token efficiency.
Tree- and graph-based paradigms offer alternative exploration. Tree-of-Thought (ToT)(Yao et al., 2023a) employs a hierarchical structure to explore multiple decision paths, though exponential growth limits deep efficiency. Graph-of-Thought (GoT)(Besta et al., 2024) extends this with flexible graph configurations, enabling path recalibration and information aggregation. Meanwhile, Skeleton-of-Thought (SoT)(Ning et al., 2023) generates high-level outlines before detailing, balancing speed and quality, while recursive methods like Self-Refine(Madaan et al., 2023) and Step-Back Prompting (Zheng et al., 2023) refine outputs iteratively for higher accuracy.
2.2. Minimal Free Resolution
MFR is a core tool in homological algebra (Fieldsteel and Nagel, 2021) and algebraic geometry (Botbol et al., 2021), used to analyze the algebraic structure of modules (Eisenbud, 2013; Rossi and Sharifan, 2009), revealing properties such as rank, symmetry, and relationships (Chen, 2011). In computational algebraic geometry, MFR analyzes singularities and invariants of algebraic varieties through syzygies (Evans and Griffith, 1981), optimizing the complexity and efficiency of Gröbner basis computations (Eisenbud et al., 2001; Stewart, 1993; Weispfenning, 1992; Capani et al., 1997).
In Topological Data Analysis (TDA), MFR accelerates persistent homology computations (Wasserman, 2018), enhancing the stability of topological analysis for high-dimensional data. In physics and bioinformatics, MFR is used to analyze Calabi-Yau manifold singularities (Polchinski, 1994), gauge field structures (Hitchin, 2003), and the modular properties of gene regulatory networks (Li et al., 2011).
The advantages of MFR in optimizing computational complexity and revealing algebraic structures provide a new perspective for its integration into the study of CoT in LLM. In CoT, structured reasoning steps are required, but often face challenges of inefficiency and redundancy. We innovatively introduce MFR into CoT by decomposing the symbolic dependencies between reasoning steps, eliminating redundant calculations, optimizing intermediate processes, and enhancing transparency. This approach opens up a more efficient and structured technical pathway for LLM applications in solving complex problems.
3. Method
3.1. Preliminaries
@@
In algebraic geometry and homological algebra, minimal free resolution is an essential tool for studying the structure and properties of modules, especially when dealing with finitely generated modules over a ring.
Key Properties: In algebraic geometry and homological algebra, a ”syzygy” refers to the relations among the relations that generate a module. A syzygy is a relation among the generators. For example, if a module is defined by generators , then a syzygy is a relation of the form:
where . These relations themselves can form a module, which is called the ”first syzygy module.” A module over a ring is an algebraic structure similar to a vector space, but the scalars come from instead of a field. A module is called free if it has a basis, i.e., for some dimension (finite or infinite). Free modules are analogous to free vector spaces in linear algebra. A module is finitely generated if there exists a finite set such that . A MFR of a finitely generated module over is an exact sequence:
Where each is free (), the matrices of the maps contain no units, and the resolution is ”minimal,” meaning the number of free generators is the smallest possible. This resolution provides a standardized way to describe the structure of , reflecting both its complexity and its homological properties. A MFR is characterized by its minimality, where each uses the smallest number of generators, and the matrices in the maps contain no units. It is unique up to isomorphism, meaning any two MFRs of are naturally isomorphic. The -th Betti number of , defined as , counts the number of free generators in and reflects the complexity of the resolution.
3.2. Minimal Free Resolution enhanced CoT
@@
As shown in Figure 3, the SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs. Each reasoning step—such as problem interpretation, task decomposition, subgoal chaining, and final inference—is abstracted as generators or relations within modules, reinterpreting CoT as algebraic objects with generative structures and latent dependencies, rather than merely a flat sequence of symbols.
![[Uncaptioned image]](extracted/6365550/Fig/modul.png)
@@
Effective problem-solving requires systematic strategies that map auxiliary conditions to actionable reasoning paths. These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises). LLMs iteratively solve subproblems along these paths, linking current conditions to new conclusions, ensuring continuity and interpretability throughout the reasoning chain.
![[Uncaptioned image]](extracted/6365550/Fig/exactness.png)
@@
After introducing the above six concepts, LLMs can be directly stimulated in a conversational manner. Figure 4 illustrates the conversational framework of these six concepts. First, complex reasoning problems are defined as a free module. Then, through the Betti number and auxiliary conditions of degrees of freedom, the complex problem is reduced and decomposed. By analyzing both the module and degrees of freedom simultaneously, the model can construct multiple mappings to solve the problem. Finally, minimality and accuracy are used to select the optimal solution.
3.3. Syzygy of Thoughts
The process of LLM reasoning for complex problems can be viewed as solving a complex reasoning problem through MFR. The entire process is formalized via a MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
@@ 3.3. Syzygy of Thoughts
where:
- is the reasoning space (LLM Reasoning Space), - are the basic units (Tokens) used in the reasoning construction, - Each is induced by the LLM, representing a single-step logical deduction in the CoT.
3.3.1. SoT Reasoning.
The Algorithm 1 and Figure 3 explain the entire process of SoT. The SoT framework begins with an initial complex problem , characterized by high-dimensional structure, intricate logical dependencies, and ambiguous constraint relationships. We treat as a unified representation of interdependent substructures, setting the stage for systematic decomposition. To analyze the inherent complexity of , we introduce Betti numbers as quantitative indicators of structural complexity. Specifically, the Betti number corresponds to the number of auxiliary conditions required for decomposition, with higher values suggesting deeper structural intricacy. These values guide the subsequent generation of auxiliary components and indicate potential optimization points.
@@
Method | Math Reasoning | Gene. | Multitask | Temporal | Log. | ||||

### v3

# Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution

@@ Abstract
Chain-of-Thought prompting enhances the reasoning of large language models by decomposing problems into sequential steps, mimicking human logic and reducing errors. However, for complex tasks with vast solution spaces and ambiguous constraints, a single reasoning chain often proves insufficient. Inspired by Minimal Free Resolutions (MFR) in commutative algebra and algebraic geometry, we propose Syzygy of Thoughts (SoT)—a new framework that extends CoT by introducing auxiliary, interrelated reasoning paths. By explicitly modeling dependencies among these paths, SoT captures deeper logical structure and enables more robust, structured problem solving. MFR decomposes a module into a sequence of free modules of minimal rank, providing a principled way to analyze complex systems. Drawing on MFR’s principle of “minimal representation, precise constraints, and layer-by-layer relation tracing,” SoT systematically decomposes an original complex problem into logically complete minimal subproblems while preserving key problem characteristics and reducing reasoning length. We evaluate SoT across diverse datasets (e.g., GSM8K, MATH) and model families (e.g., GPT-4o-mini, Qwen2.5). Results show that SoT achieves inference accuracy comparable to or exceeding mainstream CoT baselines. Moreover, by aligning the sampling process with algebraic constraints, SoT improves the scalability of inference-time computation, offering both transparent reasoning and strong performance. The code is available at https://github.com/dlMARiA/Syzygy-of-thoughts
Syzygy of Thoughts: Improving LLM CoT with the Minimal Free Resolution
Chenghao Li1, Chaoning Zhang1††thanks: Corresponding author, Yi Lu2, Jiaquan Zhang1, Qigan Sun3, Xudong Wang3, Jiwei Wei1, Guoqing Wang1, Yang Yang1, Heng Tao Shen4,1 1UESTC; 2CNU; 3KHU; 4Tongji Univ.
1 Introduction
@@
The rapid advancements in natural language processing Vaswani et al. (2017); Devlin et al. (2019); Radford et al. (2018); Brown et al. (2020) have demonstrated outstanding reasoning capabilities across various tasks Zhao et al. (2023b); Min et al. (2023). In recent years, the Chain of Thought (CoT) prompting method Wei et al. (2022) has emerged as an effective strategy for simulating human reasoning by generating intermediate reasoning steps, significantly enhancing model performance in solving complex problems. The core advantage of the CoT framework lies in its ability to decompose problems into a series of explicit and easily interpretable intermediate steps, thereby improving the transparency and explainability of the reasoning process. This explicit stepwise reasoning not only facilitates in-depth analysis and debugging of model outputs but also provides a clear logical pathway for solving complex problems.
@@
However, despite its success in many tasks, CoT faces inherent limitations with high-dimensional, nonlinear, and abstract logical problems Wang et al. (2022a); Shi et al. (2023); Jin et al. (2024). Traditional CoT methods rely on fixed or heuristic decompositions, which fail to capture essential details in multidimensional problems Wang et al. (2022a). Additionally, the lack of structured planning in intermediate steps leads to redundant computations and increased inference latency Jin et al. (2024). While improvements like multi-path exploration, voting, and hierarchical strategies have helped, they still fail to fully address logical incompleteness and poor high-dimensional decomposition. Thus, challenges in robustness and accuracy remain for complex reasoning tasks.
@@
To address these challenges, we propose a new reasoning framework—Syzygy of Thoughts (SoT). “Syzygy” is derived from the Greek word syzygia (), meaning “union” or “pairing”. This method draws inspiration from the concept of MFR in algebraic geometry and computational algebra Eisenbud (2013); Rossi and Sharifan (2009); Botbol et al. (2021), systematically decomposing and reconstructing problems by constructing minimal free module sequences. The incorporation of MFR theory ensures a more rigorous and efficient problem decomposition process while preserving the core characteristics of the problem, effectively reducing redundant computations and logical inconsistencies. Specifically, SoT retains the stepwise reasoning advantages of CoT while introducing algebraic tools such as "module", "Betti numbers", "freeness", "mapping", "exactness" and "minimality" to deeply explore and systematically express the intrinsic structure of complex problems. This approach provides a new theoretical perspective for analyzing high-dimensional and multivariable problems.
@@
The main contributions of this work:
We introduce SoT, an MFR-inspired inference-time reasoning framework that generalizes CoT from a single linear chain to interdependent, syzygy-structured reasoning paths.
We formalize the reasoning process through an MFR lens by mapping algebraic notions (Module, Betti numbers, Freeness, Mapping, Exactness, Minimality) to concrete prompting and search operations, enabling structured decomposition, logical-closure verification, and redundancy pruning.
We propose an algebraically constrained sampling/selection procedure that improves the accuracy–cost trade-off and provides stable inference under sampling randomness.
We conduct extensive experiments across nine benchmarks and multiple LLM backbones, showing that SoT consistently outperforms CoT, CoT-SC, GoT, and AoT, with lower or comparable variance, supported by cost and ablation analyses.
2 Related Work
2.1 Advanced Reasoning Techniques in LLMs
Recent advances in prompting have substantially improved the reasoning of LLMs, with Chain-of-Thought (CoT) prompting Wei et al. (2022) being a seminal example. CoT elicits step-by-step solutions but can still struggle on multidimensional or nonlinear tasks (e.g., abstract logic and advanced mathematics). To mitigate these issues, extensions such as Zero-shot-CoT Kojima et al. (2022), Self-Consistency CoT Wang et al. (2022b), and Auto-CoT Zhang et al. (2022) have been proposed, together with verification-based methods including VerifyCoT Zhao et al. (2023a) and CoF-CoT Nguyen et al. (2023). Beyond linear CoT, structured prompting further strengthens reasoning. Least-to-Most Prompting (LtM) Zhou et al. (2022) decomposes tasks into simpler subproblems, while programmatic paradigms—Program of Thought (PoT) Chen et al. (2022), Chain of Code (CoC) Li et al. (2023), and Buffer of Thought (BoT) Yang et al. (2025)—integrate discrete variables and procedural execution. Algorithm of Thought (AoT) Sel et al. (2023) then iteratively synthesizes and refines intermediate results to improve token efficiency. Search-based paradigms explore multiple solution paths. Tree-of-Thought (ToT) Yao et al. (2023) performs hierarchical branching but may suffer from exponential growth, whereas Graph-of-Thought (GoT) Besta et al. (2024) generalizes this to flexible graph structures for aggregation and recalibration. Complementary approaches such as Skeleton-of-Thought (SoT) Ning et al. (2023), Self-Refine Madaan et al. (2023), and Step-Back Prompting Zheng et al. (2023) emphasize outlining and iterative refinement to improve accuracy and stability.
2.2 Minimal Free Resolution
MFR is a fundamental tool in homological algebra Fieldsteel and Nagel (2021) and algebraic geometry Botbol et al. (2021) for characterizing module structures Eisenbud (2013); Rossi and Sharifan (2009) and deriving invariants such as rank, symmetry, and relations Chen (2011). In computational algebraic geometry, syzygy-based MFR supports the study of singularities and invariants of varieties Evans and Griffith (1981) and improves the complexity of Gröbner-basis computations Eisenbud et al. (2001); Stewart (1993); Weispfenning (1992); Capani et al. (1997). Beyond pure algebra, MFR has been used to speed up persistent-homology computation in TDA Wasserman (2018) and to analyze structures arising in physics and bioinformatics, including Calabi–Yau singularities Polchinski (1994), gauge fields Hitchin (2003), and gene regulatory networks Li et al. (2011). These strengths—exposing structured dependencies while reducing computational redundancy—motivate our integration of MFR into CoT for LLM reasoning. By decomposing symbolic dependencies among intermediate steps, we prune redundant branches, streamline intermediate computation, and improve the transparency and efficiency of reasoning, providing a more structured pathway for solving complex problems.
3 Method
3.1 Preliminaries
@@
In algebraic geometry and homological algebra, Minimal Free Resolution is a foundational tool for analyzing finitely generated modules over a ring, offering a structured way to reveal their internal dependencies and complexity.
Key Concepts. A module over a ring generalizes vector spaces by allowing scalars from . is free if it admits a basis: ; it is finitely generated if a finite set spans via .
A syzygy refers to a relation among generators, such as:
and the collection of all such relations forms the first syzygy module.
An MFR of is an exact sequence:
where each is free and the maps contain no unit elements. The resolution is minimal if the number of generators in each is smallest possible, making the structure unique up to isomorphism.
The -th Betti number, , quantifies the number of generators at each level and reflects the complexity of .
Motivation for SoT. As illustrated in Figure 3, we reinterpret CoT reasoning within LLMs through the lens of MFR. Each reasoning stage—e.g., problem understanding, decomposition, subgoal chaining, and final decision—is treated as a generator or relation within a module. This enables SoT to capture latent logical dependencies and hierarchical structure, moving beyond CoT’s linear stepwise form toward a more principled and algebraically grounded representation.
@@
The SoT reasoning paradigm employs the MFR framework to model the structure of CoT reasoning paths in LLMs. Each reasoning step—such as problem interpretation, task decomposition, subgoal chaining, and final inference—is abstracted as generators or relations within modules, reinterpreting CoT as algebraic objects with generative structures and latent dependencies, rather than merely a flat sequence of symbols.
@@
Effective problem-solving requires systematic strategies that map auxiliary conditions to actionable reasoning paths. These paths must exhibit (i) directness, minimizing redundant operations, and (ii) logical soundness, grounding each step in explicit premises. LLMs iteratively solve subproblems along these paths, linking current conditions to new conclusions, ensuring continuity and interpretability throughout the reasoning chain.
@@
After introducing the above six concepts, LLMs can be directly stimulated conversationally. Figure 6 illustrates the conversational framework of these six concepts. First, complex reasoning problems are defined as a free module. Then, through the Betti number and auxiliary conditions of degrees of freedom, the complex problem is reduced and decomposed. By analyzing both the module and degrees of freedom simultaneously, the model can construct multiple mappings to solve the problem. Finally, minimality and accuracy are used to select the optimal solution.
3.2 Syzygy of Thoughts
The process of LLM reasoning for complex problems can be viewed as solving a complex reasoning problem through MFR. The entire process is formalized via an MFR sequence, where the structured representation of the reasoning process is a set of mappings induced by the LLM, forming the following sequence:
@@ 3.2 Syzygy of Thoughts
where is the reasoning space (LLM Reasoning Space), are the basic units (Tokens) used in the reasoning construction. Each is induced by the LLM, representing a single-step logical deduction in the CoT.
3.2.1 SoT Reasoning.
Figure 3 explains the entire process of SoT. The SoT framework begins with an initial complex problem , characterized by high-dimensional structure, intricate logical dependencies, and ambiguous constraint relationships. We treat as a unified representation of interdependent substructures, setting the stage for systematic decomposition. To analyze the inherent complexity of , we introduce Betti numbers as quantitative indicators of structural complexity. Specifically, the Betti number corresponds to the number of auxiliary conditions required for decomposition, with higher values suggesting deeper structural intricacy. These values guide the subsequent generation of auxiliary components and indicate potential optimization points.
@@
| Method | Math Reasoning | Gene. | Mul. | Tem. | Log. | ||||